OpenStudy (anonymous):

Y is a Poisson distributed random variable with mean λ = 7.2. Determine P(3 ≤ Y ≤ 4 ). a) 0.0658 b) 0.2450 c) 0.5618 d) 0.0945 e) 0.1300

OpenStudy (anonymous):

I think its A correct me if i am wrong

OpenStudy (kirbykirby):

Hmm I don't get that result. Remember the probability mass function for the Poisson is: \[ P(X=k)=\frac{e^{-\lambda}\lambda^k}{k!}, \text{for }k=0,1,2,\ldots\] So, \[ P(3\le X\le4)=P(X=3)+P(X=4)\].. make you you include both x=3 and x=4 since they are both included in the inequality.

OpenStudy (anonymous):

Oh I did not realize there was formula, i was going off a graph